# Characterization of coherent errors in gate layers with robustness to Pauli noise

### v2

@@ Appendix C Addressing Pauli noise neglect during coherent noise characterization
The action of a noisy channel modeled by $\mathcal{E}_{\mathcal{P},\theta}(\rho)$ as defined in Eq. (2) can be represented in terms of the vectorized form of the density matrix of $\rho$ and the PTM of the channel. It reads ${\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)=T_{\mathcal{E}_{\mathcal{P},\theta}}\left|\rho\right)=T_{U_{\theta}}T_{\mathcal{P}}T_{U_{I}}\left|\rho\right)}$ where $T_{U_{\theta}},\,T_{\mathcal{P}}\text{ and }T_{U_{I}}$ are the transfer matrices of the coherent, Pauli and the ideal unitary channels, respectively. As stated in the main text, we discuss the case of $T_{U_{I}}$ being the identity. We moreover assume that $\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)$ and $\left|\rho\right)$ can be experimentally estimated. The estimation of the 15 parameters of $T_{U_{\theta}}$ can be written as a minimization problem over the estimator $\hat{T}_{U_{\theta}}$.
@@
With the assumption $T_{\mathcal{P}}$ and $T_{U_{\theta}}$ being close to the identity. The minimization of the function
@@
Given that the noise channel is close to the identity, we take the approximation that the diagonal matrix $T_{\mathcal{P}}$ commutes with the matrices $T_{U_{\theta}}$ and $T_{U_{\theta}^{\prime}}$. Furthermore, because of because of the unitarity of the coherent rotation channel, $T_{U_{\theta}}^{\top}{T_{U_{\theta}}}=\mathbb{I}$ is valid. Then, the minimization problem boils down to
@@
The off-diagonal part of $T_{U_{\theta}}^{\top}T_{U_{\theta^{\prime}}}$ is anti-symmetric. Consequently, $T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}$ is diagonal and smaller or equal than ${2~{}\mathbb{I}}$. Thereby, all involved matrices are diagonal. From ${1+T_{\mathcal{P}}^{2}\geq 2~{}T_{\mathcal{P}}}$, it follows that the expression in Eq. (16) is minimal when ${T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}=2~{}\mathbb{I}}$. This equation has a unique solution corresponding to $T_{U_{\theta}}={T}_{U_{\theta^{\prime}}}$, which finally implies $\theta=\theta^{\prime}$.
This statement implies that we can optimize for the coherent noise term without considering the present Pauli noise, allowing to apply the closed form solution in Eq. (12) or conduct a least squares optimization method for $\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-\hat{T}_{U_{\theta}}\left|\rho\right)\rVert^{2}$.
@@
We aim to provide a step-by-step overview of the circuit construction for gate-layer characterization. Each circuit has three main parts: state preparation, the gate layer of interest, and measurement. Since we prepare and measure the states in a Pauli basis, those two parts only require the local gates $X$, $H$, $S$, and $S^{\dagger}$. We choose a set of $L$ states to prepare, ensuring that there is no correlation in the preparation of nearby qubits. These states can be selected randomly; however, as described in Sec. III, we choose a more structured set of $216$ initial states that are all combinations of the states $\ket{+}$, $\ket{-}$, |$\ket{i}$, |$\ket{-i}$, |$\ket{1}$, and |$\ket{0}$ prepared by applying the gates $H$, $HX$, $SH$, $SHX$, $X$, respectively $\mathbb{I}$ to the state |$\ket{0}$. Next, we apply $j$ and $j+1$ iterations of the gate-layer, where $j$ is an integer. At the final measurement stage, we need a basis that allows us to reverse the action of the ideal unitary on each subsystem of connected qubits. For every neighboring pair of qubits, $q_{i}$ and $q_{j}$, the required measurement basis must encompass the subsystem that includes these two qubits along with all other qubits sharing a two-qubit gate with them. To achieve this, we use a complete Pauli measurement basis for each subsystem, which consists of all possible combinations of measuring each qubit in the $x$, $y$, or $z$ basis.
@@
    Prepare the system in the eigenstate of a random $n$-qubit Pauli operator $P$ by applying to each qubit $H$, $HX$, $SH$, $SHX$, $X$ or $\mathbb{I}$, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
@@
    For each $k$-qubit subsystem, aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$. Perform the minimization of Eq. (5).
Figure 7: Coherent noise mitigating circuit element. The circuit element comprises 4 $C\!X$ gates, 4 fixed $z$-rotations by $\pi/2$, respectively $-\pi/2$, and 15 parameterized single-qubit rotations. The first 9 rotations (blue) account for different two-qubit coherent errors, and the final 6 rotations (red) mitigate single-qubit errors. Without the first $C\!X$ gate, the circuit element represents a corrected $C\!X$ gate. We represent the element by the symbol on the right-hand side, where $\vec{\theta}$ corresponds to the angles of the 15 parameterized rotations.
@@ Appendix E Coherent noise drift
We conducted the characterization of the gate layer represented in Fig. 3 again after 6 hours as well as 19 days. The shifts in the characterization results between those three experiments are shown in Fig. 6. First, we observe that the leading coherent errors in the system did not change over the 19 days. However, as expected, the changes observed for the two experiments within 6 hours are significantly smaller than the differences over the complete time span. After 6 hours, the maximum deviation for single-qubit coherent errors was $0.009$ radians and $0.004$ radians on average. The two-qubit errors changed maximally by $0.008$ radians and on average by $0.003$ radians. In the total period (i.e. after 19 days), the most significant single-qubit coherent errors have changed by maximally $10\%$ (corresponding to about $0.03$ radians). The two-qubit coherent rotation errors have changed by maximally $0.018$ radians. The average deviation was $0.007$ radians for the single-qubit errors and $0.003$ radians for the two-qubit errors. These results suggest that the platform is affected by systematic coherent errors which are not influenced by the hourly and daily re-calibrations of the system during which qubits’ frequency and readout angle, as well as pulses’ amplitude and phase of basic single- and two-qubit gates, are calibrated [32].
@@
Figure 8: Complete data set from coherent error mitigation. Collected data of the coherent error mitigated experiment of Fig. 3. The blue curves display the unmitigated results as shown in Fig. 3 (a). Red shows the curves of the experiment with coherent mitigation according to Fig. 3 (b). Each of the 9 different panels exhibits a different initial 2-qubit state.
Appendix F Coherent noise correcting circuit element
@@
To conduct echo-experiments that allow for correcting coherent errors we require a parameterized circuit element capable of representing the characterized rotations. For a starting point of the experiment, when all coherent noise correction terms are set to zero, the circuit must correspond to the identity circuit. Single-qubit rotations are readily available in most of the leading quantum computing architectures. To invert the two-qubit rotations, an entangling gate is necessary. We use the $C\!X$ gate as it is a native gate of the chosen platform. By interleaving 15 single-qubit rotations into a structure of four $C\!X$ gates and four fixed single-qubit rotations, as shown in Fig. 7, a circuit element with 15 parameters is created. Each of the 15 parameters, which we will denote as $\vec{\theta}$ corresponds to a different single- or two-qubit rotation. The element corresponds to the identity if all parameters are set to zero.
Appendix G Complementary mitigation data
@@
Fig. 8 displays the complete data set used to create Fig. 5. The influence on the expectation values that ideally are 00 for all number of repetitions is clearly visible for all states, as they disperse much less in the mitigated case. Furthermore, the expectation values of the operators, of which the prepared state is an eigenstate, are closer to the ideal value of $1$ in the mitigated case. While the mitigation did not work equally well for all 9 states, the whole data set gives a clear indication of the positive effect of the presented coherent noise mitigation scheme.

### v3

@@ Appendix C Addressing Pauli noise neglect during coherent noise characterization
The action of a noisy channel modeled by $\mathcal{E}_{\mathcal{P},\theta}(\rho)$ as defined in Eq. (2) can be represented in terms of the vectorized form of the density matrix of $\rho$ and the PTM of the channel. It reads ${\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)=T_{\mathcal{E}_{\mathcal{P},\theta}}\left|\rho\right)=T_{U_{\theta}}T_{\mathcal{P}}T_{U_{I}}\left|\rho\right)}$, where $T_{U_{\theta}},\,T_{\mathcal{P}},\text{ and }T_{U_{I}}$ are the transfer matrices of the coherent, Pauli, and ideal unitary channels, respectively. As stated in the main text, we discuss the case of $T_{U_{I}}$ being the identity. Moreover, we assume that $\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)$ and $\left|\rho\right)$ can be experimentally estimated. The estimation of the 15 parameters of $T_{U_{\theta}}$ can be written as a minimization problem over the estimator $\hat{T}_{U_{\theta}}$.
@@
Assume $T_{\mathcal{P}}$ and $T_{U_{\theta}}$ are close to the identity. The minimization of the function
@@
Given that the noise channel is close to the identity, we take the approximation that the diagonal matrix $T_{\mathcal{P}}$ commutes with the matrices $T_{U_{\theta}}$ and $T_{U_{\theta}^{\prime}}$. Furthermore, because of the unitarity of the coherent rotation channel, $T_{U_{\theta}}^{\top}{T_{U_{\theta}}}=\mathbb{I}$ is valid. Then the minimization problem boils down to
@@
The off-diagonal part of $T_{U_{\theta}}^{\top}T_{U_{\theta^{\prime}}}$ is antisymmetric. Consequently, $T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}$ is diagonal and less than or equal to ${2~{}\mathbb{I}}$. Thus, all involved matrices are diagonal. From ${1+T_{\mathcal{P}}^{2}\geq 2~{}T_{\mathcal{P}}}$, it follows that the expression in Eq. (16) is minimal when ${T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}=2~{}\mathbb{I}}$. This equation has a unique solution corresponding to $T_{U_{\theta}}={T}_{U_{\theta^{\prime}}}$, which finally implies $\theta=\theta^{\prime}$.
This statement implies that we can optimize for the coherent noise term without considering the present Pauli noise, allowing us to apply the closed-form solution in Eq. (12) or conduct a least-squares optimization method for $\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-\hat{T}_{U_{\theta}}\left|\rho\right)\rVert^{2}$.
@@
We aim to provide a step-by-step overview of the circuit construction for gate-layer characterization. Each circuit has three main parts: state preparation, the gate layer of interest, and measurement. Since we prepare and measure the states in a Pauli basis, those two parts only require the local gates $X$, $H$, $S$, and $S^{\dagger}$. We choose a set of $L$ states to prepare, ensuring that there is no correlation in the preparation of nearby qubits. These states can be selected randomly; however, as described in Sec. III, we choose a more structured set of $216$ initial states that are all combinations of the states $\ket{+}$, $\ket{-}$, |$\ket{i}$, |$\ket{-i}$, |$\ket{1}$, and |$\ket{0}$, prepared by applying the gates $H$, $HX$, $SH$, $SHX$, $X$, and $\mathbb{I}$ respectively to the state |$\ket{0}$. Next, we apply $j$ and $j+1$ iterations of the gate-layer, where $j$ is an integer. At the final measurement stage, we need a basis that allows us to reverse the action of the ideal unitary on each subsystem of connected qubits. For every neighboring pair of qubits, $q_{i}$ and $q_{j}$, the required measurement basis must encompass the subsystem that includes these two qubits along with all other qubits sharing a two-qubit gate with them. To achieve this, we use a complete Pauli measurement basis for each subsystem, which consists of all possible combinations of measuring each qubit in the $x$, $y$, or $z$ basis.
@@
    Prepare the system in the eigenstate of a random $n$-qubit Pauli operator $P$ by applying $H$, $HX$, $SH$, $SHX$, $X$, or $\mathbb{I}$ to each qubit, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
@@
    For each $k$-qubit subsystem, aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$. Carry out the minimization of Eq. (5).
Figure 7: Coherent noise mitigating circuit element. The circuit element comprises four $C\!X$ gates, four fixed $z$-rotations by $\pi/2$, or $-\pi/2$, and 15 parameterized single-qubit rotations. The first nine rotations (blue) account for different two-qubit coherent errors, and the final six rotations (red) mitigate single-qubit errors. Without the first $C\!X$ gate, the circuit element represents a corrected $C\!X$ gate. We represent the element by the symbol on the right-hand side, where $\vec{\theta}$ corresponds to the angles of the 15 parameterized rotations.
@@ Appendix E Coherent noise drift
We conducted the characterization of the gate layer represented in Fig. 3 again after 6 hours and after 19 days. The shifts in the characterization results between those three demonstrations are shown in Fig. 6. First, we observe that the leading coherent errors in the system did not change over the 19 days. However, as expected, the changes observed for the two demonstrations within 6 hours are significantly smaller than the differences over the complete time span. After 6 hours, the maximum deviation for single-qubit coherent errors was $0.009$ radians and $0.004$ radians on average. The two-qubit errors changed maximally by $0.008$ radians and on average by $0.003$ radians. In the total period (i.e. after 19 days), the most significant single-qubit coherent errors changed by maximally $10\%$ (corresponding to about $0.03$ radians). The two-qubit coherent rotation errors changed by maximally $0.018$ radians. The average deviation was $0.007$ radians for the single-qubit errors and $0.003$ radians for the two-qubit errors. These results suggest that the platform is affected by systematic coherent errors which are not influenced by the hourly and daily recalibrations of the system during which qubits’ frequency and readout angle, as well as pulses’ amplitude and phase of basic single- and two-qubit gates, are calibrated [32].
@@
Figure 8: Complete data set from coherent error mitigation displayed partially in Fig. 4. The blue curves display the unmitigated results as shown in Fig. 4 (a). Red shows the curves of the demonstration with coherent mitigation according to Fig. 4 (b). Each of the 9 different panels exhibits a different initial two-qubit state.
Appendix F System specifications
Table 1 presents the system specifications of the IBM Quantum Falcon processor ibm_lagos as provided by IBMq [32].
Qubits Q1 Q2 Q3 Q4 Q5 Q6 Q7
readout-error [$10^{-2}$] 1.3 0.7 0.5 1.9 2.1 2.5 1.1
[…]
sx-error [$10^{-4}$] 2.8 2.3 1.9 1.9 2.7 1.9 3.0
x-error [$10^{-4}$] 2.8 2.3 1.9 1.9 2.7 1.9 3.0
CX-gate 1-2 6-7
CX-error [$10^{-3}$] $6.8$ $7.2$
gate time [$\mathrm{ns}$] $305$ $291$
Table 1: Specification of the IBM Quantum Falcon processor ibm_lagos as provided by IBMq [32] at the time of the characterization on September 22, 2022. The numbering of the qubits refers to the numbers in Fig. 3 and Fig. 6.
Appendix G Coherent noise correcting circuit element
@@
To conduct echo experiments that allow for correcting coherent errors we require a parameterized circuit element capable of representing the characterized rotations. Initially, when all coherent noise correction terms are set to zero, the circuit must correspond to the identity circuit. Single-qubit rotations are readily available in most of the leading quantum computing architectures. To invert the two-qubit rotations, an entangling gate is necessary. We use the $C\!X$ gate as it is a native gate of the chosen platform. By interleaving 15 single-qubit rotations into a structure of four $C\!X$ gates and four fixed single-qubit rotations, as shown in Fig. 7, a circuit element with 15 parameters is created. Each of the 15 parameters, which we will denote by $\vec{\theta}$ corresponds to a different single- or two-qubit rotation. The element corresponds to the identity if all parameters are set to zero.
Appendix H Complementary mitigation data
@@
Fig. 8 displays the complete data set used to create Fig. 5. The influence on the expectation values that ideally are 00 for all numbers of repetitions is clearly visible for all states, as they disperse much less in the mitigated case. Furthermore, the expectation values of the operators, of which the prepared state is an eigenstate, are closer to the ideal value of $1$ in the mitigated case. While the mitigation did not work equally well for all nine states, the whole data set gives a clear indication of the positive effect of the coherent noise mitigation scheme presented.